OpenStudy (anonymous):

Change the velocity of 40 km per hr to feet per second. with medal

OpenStudy (compphysgeek):

I dont know how many feet there are in one kilometer but what you need to do is multiply \[ \frac{40 km}{hr} \times \frac{1 hr}{3600 s} \times \frac{x ft}{1 km}\] so that the unit km and hr cancel out with each other and you only have the units left that you want

OpenStudy (anonymous):

0.011111111 then next step

OpenStudy (anonymous):

times to 3280.84?

OpenStudy (anonymous):

so the answer is 36.45

OpenStudy (compphysgeek):

yes, \( \frac{40 km}{hr} = \frac{0.01111 km}{s}\). in the second step you get rid of the km by whatever the fraction \( \frac{x ft}{km} \). how many feet are there in a km?

OpenStudy (anonymous):

3280.84

OpenStudy (compphysgeek):

then you got it right. 40 km/hr are 36.45 ft/s

OpenStudy (anonymous):

yes